### 23 Nov 2023 v1

# Shortcut Bias Mitigation via Ensemble Diversity Using Diffusion Probabilistic Models

Abstract
Spurious correlations in the data, where multiple cues are predictive of the target labels, often lead to a phenomenon known as simplicity bias, where a model relies on erroneous, easy-to-learn cues while ignoring reliable ones. In this work, we propose an ensemble diversification framework exploiting Diffusion Probabilistic Models (DPMs) for shortcut bias mitigation. We show that at particular training intervals, DPMs can generate images with novel feature combinations, even when trained on images displaying correlated input features. We leverage this crucial property to generate synthetic counterfactuals to increase model diversity via ensemble disagreement. We show that DPM-guided diversification is sufficient to remove dependence on primary shortcut cues, without a need for additional supervised signals. We further empirically quantify its efficacy on several diversification objectives, and finally show improved generalization and diversification performance on par with prior work that relies on auxiliary data collection.
1 Introduction
Deep Neural Networks (DNNs) have achieved unparalleled success in countless tasks across diverse domains. However, they are not devoid of pitfalls. One such downside manifests in the form of shortcut learning, a phenomenon whereby models latch onto simple, non-essential cues that are spuriously correlated with target labels in the training data [5, 23, 21]. Often engendered by the under-specification present in data, this simplicity bias often presents easy and viable shortcuts to performing accurate prediction at training time, regardless of a model’s alignment to the downstream task. For instance, previous work has found models to incorrectly rely on background signals for object prediction [31, 3], or to rely on non-clinically relevant metal tokens to predict patient conditions from X-Ray images [34]. The phenomenon of shortcut learning has often been found to lead to significant drops in generalization performance [1, 27, 13, 10, 34]. Leveraging shortcut cues can also be harmful when deploying models in sensitive settings. For example, shortcuts can lead to the reinforcement of harmful biases when they endorse the use of sensitive attributes such as gender or skin color [28, 32, 21].
Addressing the simplicity bias in machine learning models has been a focus of extensive research. Numerous studies have aimed to encourage models to use a broader and more diversified set of predictive cues, especially when dealing with training data that lacks explicit shortcut cue labels. A variety of these methods have been input-centric, designed to drive models to focus on different areas of the input space [26, 14], while others have focused on diversification strategies that rely on auxiliary data for disagreement [16, 9]. The latter approach, in particular, has been instrumental in developing functionally diverse models that exhibit robustness to shortcut biases, albeit with the limitation of requiring access to expensive auxiliary data.
[…]
The primary objective of this work is to mitigate shortcut learning tendencies, particularly when they result in strong, unwarranted biases, access to additional data is expensive, and different features may rely on similar areas of the input space. To achieve this objective, we propose an ensemble framework relying on unlabelled ood data for diversification. To overcome the challenges of the past, we aim to synthetically generate the data for model disagreement, and thus avoid the impracticality of ood data collection. We posit that the synthetic data should: first, lie in the manifold of the data of interest; and second, be at least in part free of the same shortcuts as the original training data. We leverage Diffusion Probabilistic Models (DPMs) to generate synthetic counterfactuals for ensemble disagreement.
Although the in-depth study of the generalization properties of diffusion sampling mechanisms is beyond the scope of this paper, we make the crucial observation that even in the presence of correlated features in the data, appropriately trained DPMs can be used to generate samples that break the shortcut signals present at training time. We show that this important characteristic arises at specific training intervals, and that it can be leveraged for shortcut cue mitigation. We hypothesize that diversification, and shortcut mitigation, can be achieved via ensemble disagreement on these DPM-generated counterfactuals, providing models with an opportunity to break the spurious correlations encouraged during training. Remarkably, without a need to control for the diversity in our synthetic samples, our experiments confirm that the extent and quality of our diffusion-guided ensemble diversification is on par with existing methods that rely on additional data.
Our contributions are three-fold:
  1. 1.
    We show that DPMs can generate feature compositions beyond data exhibiting correlated input features.
  2. 2.
    We demonstrate that ensemble disagreement is sufficient for shortcut cue mitigation.
  3. 3.
    We show that counterfactuals from appropriately trained DPMs can be used to achieve state-of-the-art diversification and shortcut bias mitigation.
2 Related Work
2.1 Overcoming the Simplicity Bias
To overcome the simplicity bias, various methods have strived to find ways to drive models to use a diversified set of predictive signals. Copious literature has explored methodologies to avoid or mitigate shortcut cue learning when labels for the shortcut cues were present in the training data [10, 7, 28, 6, 18]. As it is not always easy to predict shortcut signals, it is generally hard or impossible to have appropriate labels associated with the shortcut cues.
A common solution has historically been to train an ensemble of models, where each predictor is encouraged to use different cues for inference, thus increasing the chance for some to avert their attention from simple cues. To ensure model diversity, many methods have been proposed. For example, prior work has shown how a set of models can be trained while ensuring mutual orthogonality of the input gradients, effectively driving models to attend to different locations of the input space for prediction [17, 25, 26, 14]. These input-centric methods, however, may be at a disadvantage in cases where different features must attend to the same area of the input space.
Instead, a more direct approach has been via the direct diversification of the models’ outputs within the ensemble. This approach hinges on the availability of –unlabelled– auxiliary, and effectively Out-Of-Distribution (ood), data that is, at least in part, free of the same shortcuts as the original training data. Through a diversification objective, then, the models are made to disagree on these the ood samples, effectively fitting functions with different extrapolation behaviors [9, 16, 22]. Even in this case, the dependency on auxiliary ood data poses some limitations, as this is often not readily accessible, and can be costly to procure.
2.2 DPMs Modelling and Generalization
In recent years, Diffusion Probabilistic Models have emerged as a transformative generative tool. Numerous studies have underscored their prowess in generating synthetic images, which can then be harnessed to enrich datasets and bolster classification performance [20, 2, 33]. In several cases, DPMs have been shown to transcend the surface-level statistics of the data, making them invaluable in understanding data distributions and features [4, 33, 30]. Moreover, recent studies have indicated the ability of DPMs to achieve feature disentanglement via denoising reconstructions [8, 30, 15]. In particular, work in [29] has shown how text-guided generative models can represent disentangled concepts, and how through algebraic manipulation of their latent representations it is possible to compositionally generalize to novel and unlikely image feature combinations. In our work, we test the edge case of DPMs trained with data exhibiting correlated input features, and when suitably trained, we find them capable of generating samples with novel feature combinations.
3 Methods
The following sections describe the experiment pipeline of this paper. In §3.1, as a testbed for our experiments, we introduce an extreme data setup where the task labels are fully correlated with each of the input features. In §3.2, we introduce the datasets considered in our experiments. In §3.3 we then briefly summarize diffusion training and sampling, before presenting the diversification objectives considered in §3.4 and our experimental design in §3.5.
3.1 Wisconsin Card Sorting Test for Machine Learners
To isolate and investigate shortcut biases, we employ the Wisconsin Card Sorting Test for Machine Learners (WCST-ML), a method devised to dissect the shortcut learning behaviors of deep neural networks [21]. WCST-ML provides a systematic approach to creating datasets with multiple cues, designed to correlate with the target labels. Specifically, given $K$ cues $i_{1},i_{2},\ldots,i_{K}$, the method produces a diagonal dataset $\mathcal{D}_{\text{diag}}$ where each cue $i_{k}$ is equally useful for predicting the labels $Y$. This level playing field is instrumental in removing the influence of feature dominance and spurious correlations, thereby allowing us to observe a model’s preference for certain cues under controlled conditions. To rigorously test these preferences, WCST-ML employs the notion of off-diagonal samples. These are samples where the cues are not in a one-to-one correspondence with the labels, but instead align with only one of the features under inspection. By evaluating a model’s performance on off-diagonal samples, according to each feature, we can test and achieve an estimate of a model’s reliance on the same.
Figure 2: Training and Diffusion Generated samples for ColorDSprites and UTKFace based on WCST-ML. While training on images showcasing a correlated set of features, sampling from DPMs at appropriate fidelity levels can generate novel objects beyond the combinations of features observed during training (marked right-hand side images).
3.2 Datasets
As the experimental grounds for our study, we leverage two representative datasets, a color-augmented version of DSprites [12] and UTKFace [35], previously shown to contain features leading to strong cue learning biases [21].
DSprites: DSprites includes a comprehensive set of symbolic objects generated with variations in five latent variables: shape, scale, orientation, and X-Y position. We augment this dataset with a color dimension and remove redundant samples (e.g. due to rotations), resulting in $2,862,824$ distinct images, which we refer to as ColorDSprites. Within the context of WCST-ML, ColorDSprites permits the examination of shortcut biases in a highly controlled setup.
UTKFace: UTKFace provides a dataset of $23,708$ facial images annotated with attributes like age, gender, and ethnicity. Unlike DSprites, UTKFace presents a real-world, less controlled setup to study bias. Its inherent complexity and diversity make it an ideal candidate for understanding the model’s cue preferences when societal and ethical concerns are at stake.
Operationalizing WCST-ML Across Datasets
We follow the set-up in [21] and construct a balanced dataset $\mathcal{D}_{\text{diag}}$, which includes a balanced distribution of cues, coupled with their corresponding off-diagonal test sets (one for each feature). For both datasets, we define a balanced number of classes $L$ for each feature under investigation. Where the number of feature values exceeds $L$, we dynamically choose ranges to maintain sample balance with respect to each new feature class. For instance, for the continuous feature ‘age’ in UTKFace, we dynamically select age intervals to ensure the same $L$ number of categories as other classes, as well as sample balance within each category. We consider sets of features previously found to lead to strong simplicity biases. For ColorDSprites, we consider $K_{DS}=4$, features $\{color,orientation,scale,shape\}$, and $L=3$ as constrained by the number of shapes in the dataset. Within UTKFace we consider $K_{UTK}=3$, features $\{ethnicity,gender,age\}$, and $L=2$ as constrained by the binary classification on gender. For each dataset we create one diagonal subset of fully correlated features and labels, available at training time, and $K_{DS}$ and $K_{UTK}$ feature-specific off-diagonal datasets to serve for testing the models’ shortcut bias tendencies.
3.3 DPMs and Efficient Sampling
We utilize Diffusion Probabilistic Models (DPMs) to generate synthetic data for our experiments. DPMs operate by iteratively adding or removing noise from an initial data point $x_{0}$ through a stochastic process governed by a predefined noise schedule. The forward, reverse, and sampling processes can be summarized as follows:
[…]
where $x_{t}$ represents the data at each timestep $t$. In the forward process, noise is added to $x_{0}$ to transform it into $x_{t}$. The reverse process then aims to reconstruct $x_{t-1}$ from $x_{t}$, starting from $x_{T}\sim\mathcal{N}(0,I)$. The term $\mu_{\theta}(x_{t},t)$ is the neural network’s output, predicting the denoised data at timestep $t$, $\sigma_{t}$ is the noise standard deviation at that timestep, and $\epsilon_{t}\sim\mathcal{N}(0,I)$ is a noise term. The objective is to learn reverse transition distributions to map noisy data back to the original data $x_{0}$. To this extent, we follow the approach detailed in [19]. We parameterize the denoising model $\hat{x}_{\theta}(x_{t})$ to predict the clean data from the noisy observations, with loss:
$L_{\theta}=\frac{1+\alpha_{t}^{2}}{\sigma_{t}^{2}}\|x-\hat{x}_{\theta}(x_{t})\|^{2},$
where $\alpha_{t}$ and $\sigma_{t}$ follow a cosine schedule.
To facilitate efficient sampling, we employ Denoising Diffusion Implicit Models (DDIM) [24], a first-order ODE solver for DPMs [19, 11], utilizing a predictor-corrector scheme to minimize the number of sampling step. In particular, we set $\sigma_{t}=0$ in the predictor-corrector scheme to enhance the efficiency of the sampling process.
In our setup, we consider DPMs trained at different fidelity levels, postulating that, generally, DPMs trained for longer durations should more closely fit the data distribution at hand. Therefore, in our experiments, we use the ‘number of diffusion training epochs’ as a proxy for the DPM fidelity on the modeled distribution.
3.4 Ensemble Training and Diversification
[…]
$\mathcal{L}=\mathcal{L}_{\text{xent}}+\gamma\ \mathcal{L}^{\text{obj}}_{\text{div}}$ (1)
Where $\mathcal{L}_{\text{xent}}$ is the cross-entropy loss, $\mathcal{L}^{\text{obj}}_{\text{div}}$ is the diversification term for a particular objective, and $\gamma$ is a hyper-parameter used to modulate the importance of diversification within the optimization objective. To impart diversity to the ensemble, we investigated five diversification objectives denoted by ${{obj}\in\{\ \text{div},\ \text{cross},\ l_{1},\ l_{2},\ \text{kl}\}}$.
Figure 3: OOD sample frequency for DPMs trained at different fidelities on ColodDSprites. Three intervals are prominent: burn-in, characterized by high ood sample frequency, but poor sample quality; originative, where the model has learned the manifold of the data while still displaying capabilities for ood sample generation; and exact, characterized by near-perfect samples, but a significantly reduced ood sample generation.
We consider the $l_{1}$ and $l_{2}$ baseline objectives, designed to induce diversity by maximizing the pairwise distance between any two model outputs and the moving average of the ensemble prediction, thus:
$L^{l_{1}}_{\text{reg}}$ $=-\frac{1}{N_{m}}\sum_{i=1}^{N_{m}}\left\|f_{i}-\frac{1}{N_{m}}\sum_{j=1}^{N_{m}}f_{j}\right\|_{1}$ (2)
$L^{l_{2}}_{\text{reg}}$ $=-\frac{1}{N_{m}}\sum_{i=1}^{N_{m}}\left\|f_{i}-\frac{1}{N_{m}}\sum_{j=1}^{N_{m}}f_{j}\right\|_{2}^{2}$ (3)
The $cross$ objective diversifies the predictions by minimizing the negative mutual cross-entropy of any two models:
$\mathcal{L}^{\text{cross}}_{\text{reg}}=-\sum_{i\neq j}\frac{\text{CE}(f_{i},\text{argmax}(f_{j}))+\text{CE}(f_{j},\text{argmax}(f_{i}))}{2}$ (4)
The $kl$ objective aims at maximizing the $kl$ divergence between the output distributions of any two models:
$L^{kl}_{\text{reg}}=-\sum_{i\neq j}D_{KL}(f_{i}||f_{j})$ (5)
Finally, the $div$ diversification objective is adapted from [9], and can be summarized as:
$\mathcal{L}^{\text{div}}_{\text{reg}}=\sum_{i\neq j}$ $D_{KL}(p(f_{i},f_{j})||p(f_{i})\otimes p(f_{j}))$ (6)
$+\sum_{i}D_{KL}(p(f_{i})||p(y))$
where $y$ is the target distribution. The first term of the sum encourages diversity by minimizing the mutual information from any two predictors, and the second acts as a regularizer to prevent degenerate solutions.
(a) ColorDSprites
(b) UTKFace
Figure 4: Prediction diversity when diversifying via samples from diffusion models at different fidelities (training epochs), as compared to using real ood samples (OOD), and a baseline without diversification (BS).
3.5 Experiment Design
To investigate the objectives of this study, we apply the WCST-ML framework on ColorDSprites and UTKFace, creating training datasets of fully correlated feature-labels groups (Fig. 1). For ColorDSprites, we consider the features of ${K_{DS}=\{color,orientation,scale,shape\}}$, while in UTKFace we consider the features ${K_{UTK}=\{ethnicity,gender,age\}}$.
We train two DPMs on the diagonal fully-correlated sets of both ColorDSprites and UTKFace, including respectively 34998 and 1634 feature-correlated samples. As mentioned in §3.5, we consider the number of DPM training epochs to be a proxy of the diffusion model’s fidelity, or closeness, to the distribution of interest, where longer training generally leads to higher generative consistency of the target distribution. We generate $\approx 100$k samples from DPMs trained at varying number of epochs between 1 to 1K, to be used for ensemble diversification and analysis. We perform no post-processing or pruning of the generated samples, and instead wish to observe the innate ability of DPMs to generate samples beyond the training distribution.
For each of the ensemble experiments, we train a diverse ensemble comprising 100 ResNet-18 models on both ColorDSprites and UTKFace. We perform ensemble training separately with all diversification objectives considered, and for each objective, we perform ablation studies applying the diversification objectives to the data generated by the DPMs at different fidelity levels. For comparison, we also consider the ensemble diversification by using real ood data, randomly sampled from the off-diagonal sets, as well as a baseline without a diversification objective.
4 Results
(a) ColorDSprites
(b) UTKFace
Figure 5: Diversity comparison for ensemble trained on ood data and diffusion counterfactuals across metrics (higher is better).
4.1 Diffusion Disentanglement & OOD sampling
DPMs Exhibit Generalization Capabilities Under Correlation:
We test the ability of DPMs to transcend surface-level statistics of the data and generate samples that break the shortcut signals at training time, even when trained on correlated data. We train the DPMs on the correlated WCST-ML diagonal sample sets for both ColorDSprites and UTKFace. Fig. 2 displays the training samples for both datasets (left halves), as well as samples from the DPMs (right halves) trained for respectively 25 and 800 epochs. In the figure we observe how sampling from the trained DPMs generates previously unseen feature combinations, despite the correlated coupling of features during training (e.g. $\langle green/white,heart\rangle$, $\langle green,square\rangle$ or $\langle child,female\rangle$), suggesting a potential for innate disentangling capabilities in diffusion latent spaces. Leveraging this powerful capability is crucial in the generation of samples for disagreement, as it allows models to break from the shortcuts in the training distribution.
Early Stopping to Capture Diffusion OOD Sampling Capabilities:
To understand under which conditions sampling from DPMs can lead to the generation of samples displaying unseen feature combinations, we examine the fraction of ood samples generated by DPMs at different fidelities (§3.5). We consider ColorDSprites, given its synthetic nature, and the minimal overlap of different feature sets, making it particularly useful for this analysis. To measure the fraction of ood samples, we train a near-perfect oracle on the full ColorDSprites dataset, trained to predict the WCST-ML partitioned labels under all features. We can then consider in-distribution (id) those samples showcase fully correlated feature levels (close to the diagonal samples), and out-of-distribution (ood) those samples classified to belong to the off-diagonal WCST-ML set. Fig. 3 shows the fraction of OOD samples generated by the ColorDSprites DPM at varying levels of training fidelities. We identify at least three qualitative different intervals. An initial burn-in interval, characterized by a high frequency of ood generated samples, but which fails to capture the manifold of the data under inspection; an originative interval, where we observe a reduced number of ood samples, in favor for a generative distribution more aligned with the data to represent; and an exact interval, where the DPM’s ability to almost perfectly represent the data comes at the cost of novel emergent feature mixtures. In the context of ensemble diversification and bias mitigation, the originative interval is posed to provide the necessary information to break the simplicity shortcuts, while providing effective disagreement signals with respect to the visual cues available.
4.2 Diffusion-guided Ensemble Diversity
We trained a diverse ensemble comprising 100 ResNet-18 models on both ColorDSprites and UTKFace. The ablation studies examined the training of the ensemble with each diversification objective in turn, as well as a baseline with no imposed diversification. For both baseline and diversification experiments, we favor different training dynamics by using separate vanilla Adam optimizers for each model. We train the ensemble with a cross-entropy loss on the correlated diagonal training data, as well as an additional, objective-specific, diversification loss, computed on a separate diversification set. In our experiments, we will refer to ‘ood’ when the data for disagreement is comprised of left-out, feature-uncorrelated, samples from the original dataset; to ‘diffusion’ when the samples are generated by a DPM; and to ‘baseline’ when no diversification objective is included. We tune $\gamma$ by grid-search to provide comparable experiments under different objectives (see Suppl. §6.1)
Table 1: Results on ColorDSprites and UTKFace when using real ood samples (upper tables), or diffusion-generated counterfactuals (lower tables), for model disagreement. The feature columns report the fraction of models (not accuracy) biased towards the respective features. The final column reports the average validation accuracy for the ensemble on the left-out feature-correlated diagonal set, of the same distribution as the original training data.
OOD Disagreement
ColorDSprites
UTKFace
color orientation scale shape
valid. accuracy
(mean +/- std)
baseline 1.00 0.00 0.00 0.00 1.000 +/- 0.00
cross 0.99 0.00 0.01 0.00 0.865 +/- 0.17
div 0.86 0.01 0.11 0.02 0.818 +/- 0.22
kl 0.91 0.02 0.07 0.00 0.822 +/- 0.21
std 0.91 0.00 0.08 0.01 0.813 +/- 0.20
var 0.84 0.02 0.13 0.01 0.729 +/- 0.23
age ethnicity gender
valid. accuracy
(mean +/- std)
baseline 0.00 1.00 0.00 0.920 +/- 0.02
cross 0.01 0.71 0.28 0.865 +/- 0.04
div 0.00 0.94 0.06 0.859 +/- 0.03
kl 0.05 0.76 0.19 0.818 +/- 0.06
l1 0.02 0.62 0.36 0.847 +/- 0.07
l2 0.03 0.63 0.34 0.798 +/- 0.11
UTKFace
Diffusion Disagreement
ColorDSprites
UTKFace
color orientation scale shape
valid. accuracy
(mean +/- std)
baseline 1.00 0.00 0.00 0.00 1.000 +/- 0.00
cross 0.96 0.00 0.04 0.00 0.856 +/- 0.16
div 0.94 0.00 0.06 0.00 0.916 +/- 0.13
kl 0.89 0.01 0.10 0.00 0.786 +/- 0.20
l1 0.89 0.00 0.09 0.02 0.784 +/- 0.20
l2 0.93 0.02 0.03 0.02 0.762 +/- 0.22
age ethnicity gender
valid. accuracy
(mean +/- std)
baseline 0.00 1.00 0.00 0.920 +/- 0.02
cross 0.00 0.94 0.06 0.836 +/- 0.05
div 0.00 0.98 0.02 0.826 +/- 0.05
kl 0.00 0.94 0.06 0.837 +/- 0.06
l1 0.00 0.77 0.23 0.816 +/- 0.11
l2 0.01 0.83 0.16 0.757 +/- 0.12
UTKFace
DPM Fidelity Significantly Impacts Diversification:
We consider the case where the diversification objective is computed on samples generated by a DPM at different fidelity levels. We show in Fig. 4 the diversity achieved on the div diversification objective when training the ensemble on the same, with varying diffusion fidelity samples, and report the full set of diversification fidelity experiments in Suppl. Fig S2. In both Fig. 4a and Fig. 4b we observe the diversity of the ensemble predictions to vary significantly, from the low-diversity baseline (BS), trained without a diversification set, to the high diversity achieved with WCST-ML off-diagonal samples. For both datasets, extremely low fidelities provide little use for diversification, leading to low-diversity predictions by the ensembles. The DPM trained on ColorDSprites, however, quickly fits the synthetic data distribution and provides counterfactual samples achieving similar diversification performance to the original off-diagonal samples at early fidelity levels. Ultimately, excessive diffusion training leads to limited ood sample generation and lower diversification performance. The DPM trained on UTKFace trains longer to appropriately model the face data distribution, achieving comparable diversification levels to the ood sample set after approximately 800 training epochs. Importantly, these results highlight the benefit of appropriate DPM training and early stopping procedures to generate samples that capture the distribution at hand without displaying over-fit behaviors, affecting the potential to generate novel feature combinations and break the feature shortcuts at training time.
Diffusion-guided Diversity Leads to Comparably Diverse Ensembles:
Given the difficulty and cost of collecting auxiliary ood data for diversification, we wish to compare the level of diversification achieved with DPM counterfactual, as opposed to using real ood data. Based on our previous results in Fig. 4 (see also Suppl. Fig. S2), which do not require access to auxiliary data, we choose early-stopping DPM training at 100 and 800 epochs for ColorDSprites and UTKFace respectively, providing samples especially useful for diversification. In Fig. 5 we report the objective-wise normalized diversity achieved in each scenario by the ensemble. We find that diffusion counterfactuals can lead to almost comparable diversification performance with respect to real ood samples. In the figure, the diffusion-led diversity is almost always within 5% from the metrics achieved when using pure ood samples, both typically over 50% higher than the baseline.
(a) ColorDSprites
(b) UTKFace
Figure 6: The relationship between the change in normalized classification prediction diversity ($\Delta_{ds}$) and the change in validation accuracy of the ensemble ($\Delta_{acc}$), when trained with samples from DPMs at varying levels of fidelities. The $\Delta$s are computed with respect to baseline ensemble training, with no diversification objective. We also compare the metrics achieved by diversification with non-correlated, off-diagonal, ood data from the respective datasets.
Ensemble Diversification Breaks Simplicity Biases:
By WCST-ML, we can test each model’s bias to a cue by testing the model’s output on purposefully designed test sets [21]. We test the quality of the diversification obtained in Table 1, and ascertained the fraction of models attending to specific cues when trained on out-of-distribution and diffusion-generated samples respectively. A model was deemed to be attending to a cue if its validation accuracy on the cue-specific ood WCST-ML dataset was highest relative to all other features. Our baseline findings mirrored the observations in [21]. Specifically, each model in the baseline ensemble predominantly attended the color cue in ColorDSprites and the ethnicity cue in UTKFace, showcasing strong preferential cue bias while achieving near-perfect classification on diagonal –id– validation data. Notably, upon introducing the diversification objectives, we observed a perceptible shift in the models’ behavior, some of which averted their focus from the primary, easy-to-learn cues, turning instead to other latent cues present within the data. Among the objectives considered, $kl$, $l_{1}$, and $l_{2}$ exhibited the highest cue diversity, catalyzing the ensemble to distribute attention across multiple cues. However, this also came at the expense of a marked drop in the average ensemble performance. Conversely, the div and cross objectives yielded milder diversification, focusing primarily on the next readily discernible cues, specifically scale for ColorDSprites and gender for UTKFace, all the while maintaining a generally higher ensemble validation performance. As confirmed in Fig. 5, the diversification achieved by Diffusion counterfactuals in Table 1 is largely comparable with ood data on ColorDSprites, with respectively up to $11\%$ and $14\%$ of the models, averting their attention from the main ‘color’ cue in ColorDSprites, and respectively up to $23\%$ and $36.8\%$ of the models averting their attention to the ethnicity cue in UTKFace.
Increased Ensemble Disagreement Negatively Correlates with Ensemble id Performance
Achieving ensemble diversity by disagreement on unlabelled ood samples has previously been shown to negatively impact id ensemble performance [16] (suppl. Fig. S1), while improving ood performance on downstream tasks associated with non-shortcut cues. Generally, ensemble model selection has been an effective method to prune models that misalign with the original classification objective [9]. In this section, we wish to understand this relationship under the scope of Diffusion-guided diversification. In Fig. 6 we observe the ensemble average accuracy $\delta$ as a function of the diversity $\delta$, spanning all the diversification objectives. We highlight a few important observations. First, increased ensemble diversity via disagreement negatively correlates with average ensemble accuracy, without additional model selection mechanisms. Second, real ood data achieves the highest diversity gain, comparable to the diffusion samples at the originative stage (§4.1), $\approx[10,100]$ for ColorDSprites, and $\approx[450,800]$ for UTKFace. Finally, the diversification objectives show comparable trends, with the $div$ objective displaying marginally better diversification/accuracy performance than the others on both tasks.
5 Discussion and Conclusion
Summary. Shortcut learning is a phenomenon undermining the performance and utility of deep learning models, where easy-to-learn cues are preferentially learned for prediction, regardless of their relevance to the downstream task. We investigate the potential of DPMs to aid in the mitigation of this phenomenon by training a diverse ensemble through disagreement on synthetic counterfactuals. We show several important findings.
First, we show that DPMs can generate novel feature combinations even when trained on images displaying correlated cues, and that this phenomenon is more prevalent at particular training intervals. One significant implication of these findings is the potential of DPMs for feature disentanglement and ood sample generation, which may have interesting repercussions in several important domains including data augmentation and ood generalization.
Second, we show that diffusion-guided diversification leads models to avert attention from shortcut cues, and that diffusion counterfactuals can lead to comparable ensemble diversity without the need for expensive ood data collection.
Limitations and Future Directions. The results, and limitations, of this study open several avenues for future research. Firstly, it would be useful to understand the mechanism behind the ability of DPMs to generalize beyond the observed feature combinations, even when trained on images displaying fully correlated input features.
In our study, we find that there is a delicate balance between the level of diffusion fidelity and the effectiveness of ensemble diversification. In particular, we find it useful for a generative model to learn the manifold of the distribution under scrutiny, without (over)fitting the intricate nuances of the training data. Another interesting direction would be to delve deeper into the interplay between the fidelity of DPM-generated samples and model diversification.
A limiting factor in the disagreement objective to achieve shortcut mitigation lies in its impact on average ensemble performance. In prior work, this was partly overcome by careful model selection, but new avenues should be explored to maintain iid performance while achieving diverse extrapolation trends via disagreement.
Conclusion. This work presents a step forward in addressing the challenge of mitigating shortcut learning in DNNs. By leveraging the unique capabilities of DPMs for ensemble diversification we provide a practical method that achieves shortcut mitigation in two visual tasks, with only negligible performance loss compared to methods requiring expensive ood data collection.
Acknowledgements
Luca Scimeca acknowledges funding from Recursion Pharmaceuticals. Damien Teney is partially supported by an Amazon Research Award.
Supplementary Material
6 Supplementary Results
6.1 Calibration
We perform a hyper-parameter search to find the values of $\gamma$ for each diversification objective. We perform the search via the same methodology used in the ood main experiments, i.e. by training an ensemble of 100 ResNet-18 models on the fully correlated diagonal datasets, while diversifying a small subset of 30\$ of the original training data, randomly sampled from the de-correlated left-out set. We consider $\gamma$ values ranging from $31e-$ 3 to $11e$ 1, and monitor both the validation ensemble accuracy as well as the predictive diversity on a separate de-correlated set of validation data. Figure S1 shows the performance of each metric for different values of $\gamma$. We select the values reported in Table TABLE to be at the intersection of the accuracy and diversification trends for each model.
(a) ColorDSprites
(b) UTKFace
Figure S1: Hyperparameter search on the $\gamma$ diversification intensity for each diversification objective. For the main experiments, we select $\gamma$s at the intersection of the accuracy-diversity trends by each objective.
6.2 On the influence of DPM Fidelity to diversification
We perform experiments whereby an ensemble of 100 ResNet-18 models is trained separately with respect to all diversification objectives considered. Figure S2 shows the diversification results as a function of the fidelity of the DPM used to sample the diversification set. We observe the diversification level obtained is highly dependent on the Diffusion Fidelity level, and that appropriate early stopping procedures must be followed to achieve increased ensemble prediction diversity.
(a) cross
(b) kl
(c) L1
(d) L2
(e) ColorDSprites
(f) cross
(g) kl
(h) L1
(i) L2
(j) UTKFace
Figure S2: Ensemble diversity as enforced via samples from diffusion models trained a different fidelity levels (i.e. diffusion training epochs).
6.3 Diversification Leads to Ensemble Models Attending to Different Cues
Figure S3 illustrates a feature-centric description of 10 ensemble models trained with a diversification objective on OOD data (a) and Diffusion generated counterfactuals (b). Within each results are shown for ColorDSprites (upper panel) and UTKFace (lower panel). The variation across models is evident: several models substantially reduce their dependency on the leading cues of the respective datasets (black edges), diverging considerably from the almost identical configurations present in the baseline ensemble (red edges).
(a) OOD Disagreement
(b) Diffusion Disagreement
Figure S3: Comparison of 10 diversified models when training the ensemble while using (a) feature-uncorrelated OOD data and (b) Diffusion Samples.
6.4 On the Influence of an Increased Number of OOD Samples for Ensemble Disagreement
As per the original objective, with DPM sampling we aim to circumvent the diversification dependency on Out-Of-Distribution data, which is often not readily accessible and can be costly to procure. We test this dependency further and assess the quality of the diversification results when matching the number of OOD data used for diversification to the original training data for the ensemble. We report in Table S1 our findings. We observe the quality of the disagreement on ColorDSprites to only marginally benefit from additional disagreement samples, with approximately $10$ to $15\%$ of the models to avert their attention from the shortcut cue color. On the other hand, we observe a strong improvement in the diversification for UTKFace, mainly registered via the cross objective, where approximately $39\%$ of the models avert their attention from the ethnicity shortcut, as opposed to the original $1\%$ to $20\%$ in our previous experiments, while maintaining high predictive performance on the validation set. We speculate this gain to be due to the higher complexity of the features within the data, which may require additional specimens for appropriate diversification.
Table S1: Diversification results on ColorDSprites and UTKFace when using the same number of OOD samples as the training dataset. The feature columns report the fraction of models (in each row) biased towards the feature. The final column reports the average validation accuracy for the ensemble when tested on a left-out feature-correlated diagonal set, of the same distribution as the original training data
ColorDSprites
UTKFace
color orientation scale shape
valid. accuracy
(mean +/- std)
baseline 1.00 0.00 0.00 0.00 1.000 +/- 0.00
cross 0.92 0.00 0.08 0.00 0.849 +/- 0.16
div 0.82 0.00 0.18 0.00 0.820 +/- 0.19
kl 0.90 0.03 0.05 0.02 0.801 +/- 0.20
l1 0.90 0.01 0.07 0.02 0.799 +/- 0.21
l2 0.86 0.04 0.08 0.02 0.745 +/- 0.22
age ethnicity gender
valid. accuracy
(mean +/- std)
baseline 0.00 1.00 0.00 0.920 +/- 0.02
cross 0.00 0.73 0.27 0.858 +/- 0.04
div 0.00 0.76 0.24 0.844 +/- 0.04
kl 0.02 0.68 0.30 0.820 +/- 0.07
l1 0.00 0.66 0.34 0.832 +/- 0.09
l2 0.08 0.58 0.34 0.761 +/- 0.15
UTKFace